• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 23:53:31 on February 3,2002

    Entry number 81676

    keyword=shift summary

    started shift with access in progress to fix VDC gas flow. Apparently the gas mixing hardware is not working. Jack Segal created a "work-around".

    Simultaneously, Bogdan is directing the change of a HV crate. The right arm lower crate (used for the total absorber of the shower counter) was taken to replace a malfunctioning HV crate for the Compton setup.

    Discovered the gas problem was likely a nitrogen bottle hooked up instead of argon. Purged the delivery line in the hall and started purging chambers. VDCs taking a long time to come back.

    At end of shift, all VDCs still not showing hits over the entire chamber.